OpenStudy (anonymous):

Use implicit differentiation to find the equation of the tangent line to the curve xy^3+xy=2 at the points (1,1)

OpenStudy (anonymous):

y³ + 3xy²y' + y + xy' = 0 y' = (-y³ -y)/(3xy² + x) Equation of tangent line: y - 1 = y'(1,1) * (x-1) y-1 = -1/2 * (x-1)

OpenStudy (anonymous):

thank you can you help me with another one

OpenStudy (anonymous):

find the slope of the tangent line to the curve 3x^2-3xy-4y^3=-114 at points (2,3)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

wait actually i just got that one

OpenStudy (anonymous):

i need help with this one i got an answer but its telling me its wrong...

OpenStudy (anonymous):

6x -3y - 3xy' -12y²y' = 0 y' = (-6x + 3y)/(-3x -12y²) The value of the slope at (2,3) y'(2,3) = 1/38

OpenStudy (anonymous):

find the slope of the tangent line to the curve 1x^2+4xy+2y^3=-92 at points (-2,-4)

OpenStudy (anonymous):

2x +4y + 4xy' + 6y²y' = 0 y' = (-2x -4y)/(4x + 6y²) Value of the slope ate (-2,-4) y'(-2,-4) = 5/22

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome xD

OpenStudy (anonymous):

i have one more if you dont mind helping:)

OpenStudy (anonymous):

last one lol

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

if 4x^2+3x+xy=1 and y(1)=-6, find y'(1) by implicit dif.

OpenStudy (anonymous):

8x + 3 +y +xy' = 0 y' = (-8x -3 -y)/x at point (1,-6) y'(1) = -5

OpenStudy (anonymous):

thanks for the help:)

OpenStudy (anonymous):

any time xD